OpenStudy (anonymous):

help please Simplify: cube root of 7 over fifth root of 7 7 to the power of 1 over 5 7 to the power of 8 over 15 7 to the power of 5 over 3 7 to the power of 2 over 15

OpenStudy (anonymous):

Are you sure thats the answers?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Yup I see the answer

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

If this is what you mean\[\sqrt{7}/\sqrt[5]{7}\] then it is equal to \[7 ^{\frac{ 1 }{ 2 }}/7^{\frac{ 1 }{ 5 }} \] which is also equal with \[7^{\frac{ 1 }{ 2 }-\frac{ 1 }{ 5 }}\] and so \[7^{\frac{ 3 }{ 10 }}\] that is what i think..

OpenStudy (anonymous):

oh...what do you think it is blacklabel?

OpenStudy (anonymous):

@BlackLabel

OpenStudy (anonymous):

ow wait.. its cube root.. it is then \[7^{\frac{ 1 }{ 3 }-\frac{ 1 }{ 5 }} \] which equals \[7 ^{\frac{ 2 }{ 15}}\]

OpenStudy (anonymous):

Yup :)

OpenStudy (anonymous):

so D is correct?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

thanks! :D

OpenStudy (anonymous):

You re welcome !